# Size Winding Mechanism beyond Maximum Chaos

### v1

@@ Appendix D Exact Diagonalization Numerics
In this section, we provide the exact diagonalization results for size distribution and winding size distributions, as outlined in Eq. (S1) for $q=4$ and $N=18$. Firstly, Fig. S1 reveals that both $P(n,t)$ and $Q(n,t)$ have non-zero values only at odd operator string lengths $n$. This pattern is due to the conservation of the fermion parity. Secondly, the evolution of $P(n,t)$ and $Q(n,t)$ aligns qualitatively with the finite-$N$ results for a continuous variable of reduced size $s$, as shown in Fig.(3) of the main text. Initially, $P(n,t)$ and $|Q(n,t)|$ both peak around $n=1$, shifting to a peak at $n=1/2N$ in the later stages. Thirdly, $|Q(n,t)|$ consistently shows smaller values compared to $P(n,t)$. Early in the process, we observe notable fluctuations in both $P(n,t)$ and $|Q(n,t)|$ related to size $n$. These are not captured by the large-$q$ scramblon calculation in Fig. (3) of the main text. However, these fluctuations eventually decrease as time progresses.
@@
The winding phase depicted in Fig.S1(b) reveals a more complex structure in the context of real, finite-$N$ calculations. Initially, the early-time oscillations of the winding phase are governed by the anti-commutation properties of the fermion. These oscillations, however, are not apparent in the large-$q$ calculations that include finite-$N$ corrections. Subsequently, during the intermediate and late stages, the winding phase exhibits a linear trend across the entire operator size $n$ domain. Notably, at later times (e.g., $t=9,12$), the winding phase can be qualitatively described by the approximation formula presented in the main text Eq. (14). This relationship is approximately proportional to $12n/N-$ 1, with the slope diminishing over time. Consequently, our analysis suggests that the linear behavior of the winding phase in systems with small $N$ is predominantly influenced by finite-$N$ corrections, thereby overshadowing the large-$N$ behavior.
@@
Figure S1: We present the results of exact diagonalization for (a) $P(n,t)$ and (b) $\arg Q(n,t)$, considering a Majorana fermion system with size $N=18$, under the normalization condition $\sum_{n}P(n,t)=1$. The chosen parameters are $\Delta=\frac{1}{4}$ and a temperature that corresponds to $v=0.6$. The result is averaged over 100 random realizations. In Fig. (a), $P(n,t)$ is depicted using solid lines and solid markers, whereas $|\mathcal{Q}(s,t)|$ is represented by dashed lines and open markers. In Fig. (b), $\arg\mathcal{Q}(s,t)$ is illustrated with solid lines and open markers, which denote the winding phase, and different colors signify various real-time instances $t$. The dot-dashed line corresponds to the leading-order approximation described in the main text (Eq. (14)), applicable specifically for $t=9$ and $t=12$.

### v2

@@ Appendix B Exact Diagonalization Numerics
In this section, we provide the exact diagonalization results for size distribution and winding size distributions, as outlined in Eq. (21) for $q=4$ and $N=18$. Firstly, Fig. 5 reveals that both $P(n,t)$ and $Q(n,t)$ have non-zero values only at odd operator string lengths $n$. This pattern is due to the conservation of the fermion parity. Secondly, the evolution of $P(n,t)$ and $Q(n,t)$ aligns qualitatively with the finite-$N$ results for a continuous variable of reduced size $s$, as shown in Fig.(3) of the main text. Initially, $P(n,t)$ and $|Q(n,t)|$ both peak around $n=1$, shifting to a peak at $n=1/2N$ in the later stages. Thirdly, $|Q(n,t)|$ consistently shows smaller values compared to $P(n,t)$. Early in the process, we observe notable fluctuations in both $P(n,t)$ and $|Q(n,t)|$ related to size $n$. These are not captured by the large-$q$ scramblon calculation in Fig. 3 of the main text. However, these fluctuations eventually decrease as time progresses.
@@
The winding phase depicted in Fig. 5(b) reveals a more complex structure in the context of real, finite-$N$ calculations. Initially, the early-time oscillations of the winding phase are governed by the anti-commutation properties of the fermion. These oscillations, however, are not apparent in the large-$q$ calculations that include finite-$N$ corrections. Subsequently, during the intermediate and late stages, the winding phase exhibits a linear trend across the entire operator size $n$ domain. Notably, at later times (e.g., $t=9,12$), the winding phase can be qualitatively described by the approximation formula presented in the main text Eq. (20). This relationship is approximately proportional to $12n/N-$ 1, with the slope diminishing over time. Consequently, our analysis suggests that the linear behavior of the winding phase in systems with small $N$ is predominantly influenced by finite-$N$ corrections, thereby overshadowing the large-$N$ behavior.
@@
Figure 5: We present the results of exact diagonalization for (a) $P(n,t)$ and (b) $\arg Q(n,t)$, considering a Majorana fermion system with size $N=18$, under the normalization condition $\sum_{n}P(n,t)=1$. The chosen parameters are $\Delta=\frac{1}{4}$ and a temperature that corresponds to $v=0.6$. The result is averaged over 100 random realizations. In Fig. (a), $P(n,t)$ is depicted using solid lines and solid markers, whereas $|\mathcal{Q}(s,t)|$ is represented by dashed lines and open markers. In Fig. (b), $\arg\mathcal{Q}(s,t)$ is illustrated with solid lines and open markers, which denote the winding phase, and different colors signify various real-time instances $t$. The dot-dashed line corresponds to the leading-order approximation described in the main text (Eq. (20)), applicable specifically for $t=9$ and $t=12$.